Assume that impure copper contains only Fe, Au, Ag as impurities. After passing of 140 ampere for 482.5s, the mass of anode decreased by 22.260g and the cathode increased in mass by 22.011g. Calculate the percentage of iron and percentage of copper originally present.

Grade:11

1 Answers

Sunil Kumar FP
9 years ago
Cu ---Cu+2 +2e-
Fe ----Fe+2 +2e-
gain in weight at cathode is due to deposition of Cu =22.011g
Mole of Cu deposited at cathode =22.011/63.5 =.3466
equivalent of Cu and Fe dissolved at anode =it/96500
=140*482.5/96500 =.7
moles of Cu and Fe dissolved at anode =.70/2 =.35
mole of Fe dissolved at anode =.3500-.3466 =.0034
weight of Fe dissolved at anode =.0034*56 =.190gm
thus anode weigt loss of 220.260g contains 22.011g,.190gFe and Au+Ag =22.26-.190 =.059g
%Cu=22.011/22.26 *100 =98.88%
%Fe =.190/22.26*100 =.85%